# PCA-Relax: Deep-learning-based groupwise registration for motion correction of cardiac $T_{1}$ mapping

### v1

@@ Abstract
Quantitative MRI (qMRI) is an increasingly important tool for clinical assessment of cardiovascular diseases. Quantitative maps are derived by fitting a known signal model to a series of baseline images, while the quality of the map can be deteriorated by involuntary respiratory and cardiac motion. To correct motion, a template image is often needed to register all baseline images, but the choice of template is nontrivial, leading to inconsistent performance sensitive to image contrast. In this work, we propose a novel deep-learning-based groupwise registration framework, which omits the need for a template, and registers all baseline images simultaneously. We design two groupwise losses for this registration framework: the first is a linear principal component analysis (PCA) loss that enforces alignment of baseline images irrespective of the intensity variation, and the second is an auxiliary relaxometry loss that enforces adherence of intensity profile to the signal model. We extensively evaluated our method, termed “PCA-Relax”, and other baseline methods on an in-house cardiac MRI dataset including both pre- and post-contrast $T_{1}$ sequences. All methods were evaluated under three distinct training-and-evaluation strategies, namely, standard, one-shot, and test-time-adaptation. The proposed PCA-Relax showed further improved performance of registration and mapping over well-established baselines. The proposed groupwise framework is generic and can be adapted to applications involving multiple images.
@@ 1 Introduction
Quantitative cardiac MRI, such as $T_{1}$ and $T_{2}$ mapping [17, 19], has become an increasingly important imaging modality for non-invasive examination of heart diseases [6]. In principle, the quantitative values of $T_{1}$ or $T_{2}$ relaxometry are inferred by fitting a known parametric model to a series of baseline images with different intensities and contrasts, assuming anatomical alignment. In practice, however, the alignment assumption is often violated by the involuntary respiratory and cardiac motion of patients [26], resulting in deteriorated accuracy and precision of the quantitative mapping [9, 10]. This makes motion correction an essential post-processing step for qMRI [15, 23].
@@ 4 Conclusion
In this work, we proposed a novel template-free, deep-learning-based, groupwise registration framework, to tackle the motion correction problem for cardiac $T_{1}$ mapping. Two groupwise losses were proposed and validated: a sequence-agnostic PCA loss and a sequence-specific relaxometry loss. We extensively evaluated the proposed method, PCA-Relax, with diverse training-and-evaluation strategies on an in-house cardiac $T_{1}$ qMRI dataset. The proposed method demonstrated improved performance of registration and mapping over well-established baselines. The generic formulation of our groupwise framework allows easy extension to applications that involve multiple image registration.

### v2

# Deep-learning-based groupwise registration for motion correction of cardiac $T_{1}$ mapping

@@ Abstract
Quantitative $T_{1}$ mapping by MRI is an increasingly important tool for clinical assessment of cardiovascular diseases. The cardiac $T_{1}$ map is derived by fitting a known signal model to a series of baseline images, while the quality of this map can be deteriorated by involuntary respiratory and cardiac motion. To correct motion, a template image is often needed to register all baseline images, but the choice of template is nontrivial, leading to inconsistent performance sensitive to image contrast. In this work, we propose a novel deep-learning-based groupwise registration framework, which omits the need for a template, and registers all baseline images simultaneously. We design two groupwise losses for this registration framework: the first is a linear principal component analysis (PCA) loss that enforces alignment of baseline images irrespective of the intensity variation, and the second is an auxiliary relaxometry loss that enforces adherence of intensity profile to the signal model. We extensively evaluated our method, termed “PCA-Relax”, and other baseline methods on an in-house cardiac MRI dataset including both pre- and post-contrast $T_{1}$ sequences. All methods were evaluated under three distinct training-and-evaluation strategies, namely, standard, one-shot, and test-time-adaptation. The proposed PCA-Relax showed further improved performance of registration and mapping over well-established baselines. The proposed groupwise framework is generic and can be adapted to applications involving multiple images.
@@ 1 Introduction
Quantitative MRI (qMRI) of heart, such as $T_{1}$ and $T_{2}$ mapping [17, 19], has become an increasingly important imaging modality for non-invasive examination of heart diseases [6]. In principle, the quantitative values of $T_{1}$ or $T_{2}$ relaxometry are inferred by fitting a known parametric model to a series of baseline images with different intensities and contrasts, assuming anatomical alignment. In practice, however, the alignment assumption is often violated by the involuntary respiratory and cardiac motion of patients [26], resulting in deteriorated accuracy and precision of the quantitative mapping [9, 10]. This makes motion correction an essential post-processing step for qMRI [15, 23].
@@ 4 Conclusion
In this work, we proposed a novel template-free, deep-learning-based, groupwise registration framework, to tackle the motion correction problem for cardiac $T_{1}$ mapping. Two groupwise losses were proposed and validated: a sequence-agnostic PCA loss and a sequence-specific relaxometry loss. We extensively evaluated the proposed method, PCA-Relax, with diverse training-and-evaluation strategies on an in-house cardiac $T_{1}$ mapping dataset. The proposed method demonstrated improved performance of registration and mapping over well-established baselines. The generic formulation of our groupwise framework allows easy extension to applications that involve multiple image registration.